# FEA-Bench: A Benchmark for Evaluating Repository-Level Code Generation for Feature Implementation

### v1

@@ 1 Introduction
The remarkable text generation capabilities of large language models (LLMs) Achiam et al. (2023) have extended their impact into the domain of code generation Xu et al. (2022), which has led to the emergence of developer assistants such as Copilot, Cursor, Devin, etc. An important research topic to evaluate the effectiveness of LLMs in generating code across diverse scenarios. Many of the existing benchmarks focus on evaluating standalone programming problems, such as HumanEval, MBPP, and LiveCodeBench Chen et al. (2021); Austin et al. (2021); Jain et al. (2024). These benchmarks offer little insight into the challenges developers face in real-world projects, where codebases are composed of multiple interconnected files. In such projects, modifications in one part of the code often necessitate corresponding edits elsewhere. This type of collaborative, large-scale development is referred to as repository-level code development.
@@ 2.1 Code Large Language Models
Large language models (LLMs) have revolutionized software engineering by enabling code generation, debugging, and translation capabilities Pan et al. (2024); Li et al. (2023a); Joshi et al. (2023); Shi et al. (2024). Large-scale pre-trained LLMs such as GPT-4 Achiam et al. (2023), CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024); Zhu et al. (2024), and Qwen2.5-Coder Hui et al. (2024) have demonstrated proficiency in generating functional code across multiple programming languages. Recent advancements also include instruction-tuned models like Starcoder Li et al. (2023b), WizardCoder Luo et al. (2023), WaveCoder Yu et al. (2024b), Magicoder Wei et al. (2024) and EpiCoder Wang et al. (2025). With these advancements, code LLMs are poised to further revolutionize how developers interact with code, promising increased efficiency in software creation. Additionally, the integration of agents has further enhanced the performance of LLMs in software engineering tasks Yang et al. (2024b); Xia et al. (2024); Zhang et al. (2024). In this paper, we further evaluate the performance of current LLMs in the incremental code development scenarios at the repository level. This investigation aims to drive the research on code LLMs toward addressing more intricate software engineering challenges, thereby advancing the capabilities of these models in handling sophisticated development tasks.
@@ 3.2 Repository Collection
To determine the scope of GitHub repositories for data collection, we initially focus on the GitHub repositories corresponding to Python packages listed on the Top PyPI website  [^2] https://hugovk.github.io/top-pypi-packages/. Packages appearing on this list are generally influential Python software with high data quality, and most repositories use a unified pytest format for testing, which facilitates later execution. This leaderboard contains 8,000 Python packages. We filter out repositories that have a license and more than 1,000 pull requests, resulting in approximately 600 repositories meeting these criteria.
@@ 3.3 Task Collection and Filtering
Based on existing repository data, we crawl all pull requests and filter out those that include changes to test files as valid task instances. Given our focus on the incremental feature development task, we introduce the following steps to filter out final task instances for FEA-Bench:
Extraction of new components For each task instance candidate, we perform parsing on all Python scripts involved in the gold patch. We compare the state before and after applying the patch to identify newly added namespaces, including classes and functions, and extract their signatures and docstrings as metadata.
@@ 3.4 Benchmark Characteristics
Lite version We have also curated a subset to serve as a lite version of our dataset. This subset was filtered based on criteria including higher quality and lower difficulty. This lite version is particularly useful for evaluating systems that are computationally intensive and time-consuming. Detailed information can be found in Section A.3.
@@ 7 Conclusion
Our constructed data and experiments have certain limitations. Firstly, the quantity of high-quality data suitable for repository-level incremental development is limited. High-quality and usable pull requests for new feature development are relatively scarce. Many repository-level code developments for implementing new functionalities were committed during the early stages of repositories, without going through the rigorous code review process typical of the open-source community, resulting in lower data quality that cannot be utilized. Furthermore, the software’s early-stage developments might not even have been conducted using the GitHub platform, posing a challenge for data collection and utilization. Consequently, FEA-Bench, which is built on publicly available data and subjected to stringent filtering, may exhibit certain scenario limitations.
Secondly, due to the long context involved in repository-level code development, the cost of conducting experiments using LLMs is relatively high. Therefore, the experimental results are based on a single round generation, akin to Pass@1, which may introduce a certain level of bias into the results. Additionally, given the scarcity of API resources for models like DeepSeek-V3 and R1, some results in the main experiments presented in Table 2 are missing. We hope that more affordable models similar to DeepSeek can be further developed to facilitate research and applications in the field of repository-level code development.

### v2

@@ 1 Introduction
The remarkable text generation capabilities of large language models (LLMs) Achiam et al. (2023) have extended their impact into the domain of code generation Xu et al. (2022), which has led to the emergence of developer assistants such as Copilot, Cursor, Devin, etc. An important research topic is evaluating the effectiveness of LLMs in generating code across diverse scenarios. Many of the existing benchmarks focus on evaluating standalone programming problems, such as HumanEval, MBPP, and LiveCodeBench Chen et al. (2021); Austin et al. (2021); Jain et al. (2024). These benchmarks offer little insight into the challenges developers face in real-world projects, where codebases are composed of multiple interconnected files. In such projects, modifications in one part of the code often necessitate corresponding edits elsewhere. This type of collaborative, large-scale development is referred to as repository-level code development.
@@ 2.1 Code Large Language Models
Large language models (LLMs) have revolutionized software engineering by enabling code generation, debugging, and translation capabilities Pan et al. (2024); Li et al. (2023a); Joshi et al. (2023); Shi et al. (2024). Large-scale pre-trained LLMs such as GPT-4 Achiam et al. (2023), CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024); Zhu et al. (2024), and Qwen2.5-Coder Hui et al. (2024) have demonstrated proficiency in generating functional code across multiple programming languages. Recent advancements also include instruction-tuned models like Starcoder Li et al. (2023b), WizardCoder Luo et al. (2023), WaveCoder Yu et al. (2024b), Magicoder Wei et al. (2024) and EpiCoder Wang et al. (2025). With these advancements, code LLMs are poised to further revolutionize how developers interact with code, promising increased efficiency in software creation. Additionally, the integration of agents Luo et al. (2025) has further enhanced the performance of LLMs in software engineering tasks Yang et al. (2024b); Xia et al. (2024); Zhang et al. (2024). In this paper, we further evaluate the performance of current LLMs in the incremental code development scenarios at the repository level. This investigation aims to drive the research on code LLMs toward addressing more intricate software engineering challenges, thereby advancing the capabilities of these models in handling sophisticated development tasks.
@@ 3.2 Repository Collection
To determine the scope of GitHub repositories for data collection, we initially focus on the GitHub repositories corresponding to Python packages listed on the Top PyPI website [^2] https://hugovk.github.io/top-pypi-packages/. Packages appearing on this list are generally influential Python software with high data quality, and most repositories use a unified pytest format for testing, which facilitates later execution. This leaderboard contains 8,000 Python packages. We obtain approximately 600 repositories that meet the criteria of having a license and more than 1,000 pull requests by applying a filtering process.
@@ 3.3 Task Collection and Filtering
Based on existing repository data, we crawl all pull requests and consider those that include changes to test files as possible task instances. Given our focus on the incremental feature development task, we introduce the following steps to obtain the final task instances for FEA-Bench through filtering:
Extraction of new components For each task instance candidate, we perform parsing on all Python scripts involved in the gold patch. We compare the state before and after applying the patch to identify newly added components, including classes and functions, and extract their signatures and docstrings as metadata.
@@ 3.4 Benchmark Characteristics
Lite version We have also curated a subset to serve as a lite version of our dataset. This subset is filtered based on criteria including higher quality and lower difficulty. This lite version is particularly useful for evaluating systems that are computationally intensive and time-consuming. Detailed information can be found in Section A.3.
@@ 6 Discussion
Agentless Agentless-Lite Oracle BM25 (27K)
%Reso. %Apply %Reso. %Apply %Reso. %Apply %Reso. %Apply
DeepSeek-V3 - - 11.0 71.0 14.5 70.5 13.0 69.0
GPT-4 - - 3.5 94.5 6.0 54.0 2.0 37.0
GPT-4o 9.0 87.5 9.5 96.0 5.0 60.0 4.0 53.5
o1-mini 4.5 69.5 4.5 81.5 2.0 32.5 1.0 23.5
o1 14.0 90.5 10.0 89.5 10.0 49.5 5.0 33.5
Table 5: The performance of Agentless on FEA-Bench lite, compared to the direct retrieval by Oracle and BM25. "%Reso." refers to the resolved ratio of task instances, while "%Apply" indicates the success rate of applying the generated code edits to the repository.
6.1 Retrieved Files in Context
6.2 Output Format of Edits
@@
6.3 Evaluation of Agent Frameworks
To further evaluate the performance of current state-of-the-art methods on FEA-Bench, we conduct experiments on the lite subset using the cost-efficient Agentless and Agentless-lite frameworks Xia et al. (2024). We compare the final performance of using these agents with that of directly constructing the context using Oracle and BM25 retrieval methods, as shown in Table 5. Notably, Agentless improves the resolved ratio of task instances over BM25 for most models, with the largest gains observed for GPT-4o, o1-mini, and o1. This improvement strongly correlates with the increased success rate of applying code edits, highlighting the challenge of adhering to the required code editing format, as discussed in Section 6.2. Overall, Agentless and Agentless-lite perform similarly, with no significant advantage over the results under the Oracle setting. This indicates that current methods still have substantial room for improvement in repository-level new feature implementation.
@@ 7 Conclusion
Our constructed data and experiments have certain limitations. First, our benchmark includes only Python repositories, as Python projects are easier to execute and generally follow consistent testing frameworks. Second, the quantity of high-quality data suitable for repository-level incremental development is limited. High-quality and usable pull requests for new feature development are relatively scarce. Many repository-level code developments for implementing new functionalities were committed during the early stages of repositories, without going through the rigorous code review process typical of the open-source community, resulting in lower data quality that cannot be utilized. Furthermore, the software’s early-stage developments might not even have been conducted using the GitHub platform, posing a challenge for data collection and utilization. Consequently, FEA-Bench, which is built on publicly available data and subjected to stringent filtering, may exhibit certain scenario limitations.
Due to the long context involved in repository-level code development, the cost of conducting experiments using LLMs is relatively high. Therefore, the experimental results are based on a single round generation, akin to Pass@1, which may introduce a certain level of bias into the results. Additionally, given the scarcity of API resources for models like DeepSeek-V3 and R1, some results in the main experiments presented in Table 2 are missing. We hope that more affordable models similar to DeepSeek can be further developed to facilitate research and applications in the field of repository-level code development.
@@
Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0116308) and National Natural Science Foundation of China (62036001) . The corresponding author is Xin Zhang and Houfeng Wang.
@@ A.3 FEA-Bench lite
Beyond the lite subset, we also expect to collaborate with professional software engineers to annotate a verified subset, similar to SWE-bench Verified [^6] https://openai.com/index/introducing-swe-bench-verified/.